An equation of straight line passing through the point 2,-1 and making angle 45 ° with the 6x+5y-1=0 is?

Grade:12th pass

1 Answers

Vikas TU
6 years ago
Dear Student,
slope of line = 11 or -1/11
       (Y+1)/X-2 = 11
       Or (y+1)/x-2 = -1/11
 
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)